OpenStudy (anonymous):

1.)If I know the lengths of two sides of a right triangle, how do I find the third? 2.)Could I find the two missing side lengths of a right triangle if I only know one side length and one angle measure (other than the 90 degree angle)? 3.)Could I find the two missing angle measures if I know some of the side lengths of a right triangle? 4.)What makes a triangle a “special” right triangle? How can special right triangles help me find side lengths?

OpenStudy (anonymous):

1. a^2+b^2=c^2

OpenStudy (anonymous):

1)Pythagorean Theorem says a^2+b^2=c^2 2)Yes by using sine and cosine functions 3)Same as number 2. sine and cosine functions work both ways 4)I dont know this one sorry

OpenStudy (anonymous):

Ohh Okay. thank you :)